### v1

# Do Visual-Language Maps Capture Latent Semantics?

Abstract
Visual-language models (VLMs) have recently been introduced in robotic mapping by using the latent representations, i.e., embeddings, of the VLMs to represent the natural language semantics in the map. The main benefit is moving beyond a small set of human-created labels toward open-vocabulary scene understanding. While there is anecdotal evidence that maps built this way support downstream tasks, such as navigation, rigorous analysis of the quality of the maps using these embeddings is lacking. We investigate two critical properties of map quality: queryability and consistency. The evaluation of queryability addresses the ability to retrieve information from the embeddings. We investigate two aspects of consistency: intra-map consistency and inter-map consistency. Intra-map consistency captures the ability of the embeddings to represent abstract semantic classes, and inter-map consistency captures the generalization properties of the representation. In this paper, we propose a way to analyze the quality of maps created using VLMs, which forms an open-source benchmark to be used when proposing new open-vocabulary map representations. We demonstrate the benchmark by evaluating the maps created by two state-of-the-art methods, VLMaps and OpenScene, using two encoders, LSeg and OpenSeg, using real-world data from the Matterport3D data set. We find that OpenScene outperforms VLMaps with both encoders, and LSeg outperforms OpenSeg with both methods.
I Introduction
Mobile robots must understand the geometry and semantics of the environment to accomplish complex tasks. The ability of maps to capture semantics has increased with the development of computer vision systems, from detecting single objects to creating pixel-wise semantic segmentation of images, yielding dense semantic maps with labels in each map cell [1]. However, most semantic segmentation methods are trained to segment the image into a small pre-selected set of categories. This imposes challenges in real-world settings when a robot operates in an environment that the categories do not fully describe [2, 3].
Visual-Language Models (VLMs) are networks that jointly train a visual and language encoder to learn a mapping from text and image inputs into a common visual-language latent space. Modern, efficient transformer architectures [4], such as CLIP [5], are trained with large data sets, which enables these models to effectively have an open vocabulary, meaning visual semantics can be matched to natural language instead of a set of selected symbols. Therefore, the exploitation of VLMs in robotic mapping could allow richer semantics to be represented and to overcome the difficulty of adapting to new environments. While this work focuses on maps that encode semantics using VLMs, mapping methods have been proposed that can incorporate other modalities as well, such as audio [6]. We will collectively refer to all these maps as latent semantic maps.
Figure 1: Maps built from visual-language models have the ability to represent complex semantics, encompassing both the class of objects and their properties. We propose a benchmark for these type of maps and evaluate the quality of different state-of-the-art methods according to the consistency of their representation and their ability to be queried.
The semantic information stored in the map must be able to be queried. While complex queries have been long possible in semantic maps by inference using object ontologies [7], semantic labels support only binary comparison as they are enumerations with no notion of distance. The promise of latent semantic maps is that queries can be evaluated by directly comparing their embeddings to the embeddings of the map with a similarity metric, yielding virtually unrestricted query vocabulary. Several successful algorithms have been built on VLMs that can perform mobile robot tasks, such as navigation [8, 9, 10]. However, the analysis from the robotic mapping perspective is left anecdotal.
In this work, we propose a way to evaluate the quality of latent semantic maps. We consider that three aspects of the map capture its quality: queryability, consistency, and the breadth of the vocabulary. We concentrate on evaluating queryability and consistency because, to the best of our knowledge, no data sets exist to evaluate robotic maps that adequately capture open vocabulary semantics. Furthermore, we evaluate the consistency of the representation, i.e., the embeddings themselves.
Understanding the possibilities and limitations of the representation is essential for bringing the methods into real-world applications. Having comparable metrics for the quality of the latent semantic maps is vital for measuring the progress in this new research field, as this allows for evaluating new methods against the state-of-the-art. For this reason, we provide an open-source benchmark [^1] The source code will be made publicly available upon acceptance of this work. against which any new open vocabulary semantic representation can be tested. To demonstrate our benchmark, we evaluate two state-of-the-art methods: VLMaps [8] and OpenScene [11], using the Matterport3D [12] data set.
The main contributions of this paper are:
  1. i)
    We define queryability and consistency of representation as measures of mapping quality for latent semantic maps and propose metrics to quantitatively evaluate these properties;
  2. ii)
    We provide an open-source benchmark software for latent semantic maps, enabling evaluation of the progress in this new field of research;
  3. iii)
    We provide a comprehensive analysis of two state-of-the-art methods using the proposed benchmark.
II Related work
II-A Maps created with visual-language embeddings
Creating maps using embeddings of VLMs, especially CLIP [5], has recently gained attention, with the proposal of methods such as VLMaps [8], OpenScene [11], NLMaps [13], and Uni-Fusion [14].
VLMaps and OpenScene create a dense grid map, where each cell is associated with a single VLM embedding. The embeddings are created from RGBD images with a language-driven semantic segmentation encoder, such as LSeg [15] or OpenSeg [16], both based on CLIP, and back-projected to the 3D map. Each cell is represented as the mean of the embeddings projected to the cell. In the original work, the 3D map in VLMaps is projected to 2D by averaging the embeddings in the $z$-axis. Additionally, OpenScene learns an encoder that directly produces CLIP embeddings from the 3D point cloud. Combining these two approaches, they propose joint 2D-3D ”ensemble” features as the final representations in the map.
Unlike VLMaps and OpenScene, NLMaps creates a feature-based map, where each feature is associated with an VLM embedding. Uni-Fusion proposes a general method to create continuous maps from any perceptual data modality, including VLM embeddings. In this paper, we focus on VLMaps and OpenScene because the NLMaps and Uni-Fusion source code were not made fully publicly available at the time of writing.
II-B Open-vocabulary queries
Open-vocabulary methods aim to move beyond predicting a constrained, i.e., closed, set of labels. This means that not only should the open-vocabulary systems be able to capture synonyms and closely related terms but also attributes and related actions of the objects. This has been explored early on by formalizing the relation of labels into conceptual graphs [17], similar to what is used in semantic mapping, predicting attributes of objects [18] to the prediction of affordances and activities related to objects [11].
There are closely related fields to open-vocabulary detection and segmentation, especially open-set and zero-shot learning. In open-set learning, the task is to classify the known classes seen in training and to classify unseen classes as unseen [3]. Zero-shot learning aims to be able to classify classes that did not appear in the training data [19]. This, however, does not imply that open vocabulary is in use. Additionally, many open-vocabulary methods [20, 21, 22], even when trained with open-vocabulary methods, are evaluated in a closed-vocabulary but zero-shot setting. While this fails to capture the full potential of the open vocabulary, the results are comparable to semantic segmentation methods, where currently, fully supervised methods outperform the visual-language methods [11].
In contrast, in this work, we evaluate queryability such that the evaluation is not restricted to semantic segmentation of the map, which imposes that the map must be partitioned. Instead, each query is evaluated separately.
II-C Analysis of maps
When assessing the quality of robotic maps, most existing works concentrate on evaluating geometry using methods ranging from simple feature descriptors [23], frequency-based methods [24], to probabilistic measures [25].
Evaluating the geometric and semantic properties jointly is only relevant when the geometry and semantics are jointly estimated, e.g., in semantic Simulatenous Localization and Mapping (SLAM). Often, the metrics are qualitative, but quantitative metrics such as the localization accuracy or map reconstruction quality compared to ground truth objects are used as well [26, 27, 28]. Like many semantic mapping works, [29, 30, 31], we focus on the evaluation of the semantic properties separately, using binary classification metrics.
In VLMaps, as with other zero-shot navigation methods [9, 10], the map quality is not directly assessed, but instead, they use the success of the downstream navigation task as the metric, which does not fully evaluate the capabilities of the representation.
Because the metrics used in evaluating latent semantic maps in each prior work are different, comparing the quality of their maps is non-trivial, which is the problem we address in this paper.
III Problem statement
We aim to create a benchmark to evaluate the quality of latent semantic maps. We only consider methods where the estimation of geometry and semantics are disjoint, so we concentrate on evaluating the semantic quality of the maps. We believe that three aspects of the map capture its semantic quality: queryability, consistency, and the breadth of the vocabulary. We concentrate on evaluating queryability and consistency because, to the best of our knowledge, no data sets exist to evaluate robotic maps that adequately capture open-vocabulary semantics.
The benchmark must evaluate queryability, as it is the primary way to retrieve information from the representation and, therefore, acts as the measure of two things: first, indirectly, the ability of the representation to contain relevant information, and second, the accessibility of that information. It also must address consistency, as consistent performance within and between measuring runs, times, environments, and sensors is desirable.
IV Methods
IV-A Evaluating queryability
Because the environment consists both of background regions (e.g., floor and ceiling) and discrete object instances (e.g., a chair or a table), we propose evaluating queryability in two ways: by measuring the voxel-based coverage of the query results and the instance retrieval capability.
IV-A1 Voxel-based queryability
The first method evaluates the overall matching between the query results compared to the ground truth in a binary classification setting. This property cannot be evaluated using standard multi-class classification, as queries do not form a partition of the map. Each query produces a query result, a binary mask, over the whole map, and the same voxel might be a match of multiple queries. This is evident from the open-vocabulary perspective; e.g., a sofa might match queries ”sofa”, ”couch”, ”place to sit”, and ”soft”.
The method consists of the following steps:
  1. i)
    A map $m$ is created from each sequence from the data set, forming the set $\mathcal{M}$.
  2. ii)
    Each map $m\in\mathcal{M}$ is queried with each query $q$ in the set of queries $\mathcal{Q}$. The query result of a single query is a binary segmentation mask of the map, $\hat{\textbf{y}}_{q}=\{\hat{y}_{1},\dots,\hat{y}_{N}\}$, consisting of $N$ voxel masks $\hat{y}\in\{true,false\}$. Combined, the query results form the set of voxel-based predictions $\hat{Y}_{V}=\{\hat{\textbf{y}}_{q}\forall q\in\mathcal{Q}\}$.
  3. iii)
    For each query, the true mask $\textbf{y}_{q}$ is created from the ground truth map $m_{g}$ such, that the voxels answering query $q$ are $true$, others $false$. The true masks form a set of true labels $Y_{V}=\{\textbf{y}_{q}\forall q\in\mathcal{Q}\}$.
  4. iv)
    The binary classification metrics are calculated between $Y_{V}$ and $\hat{Y}_{V}$.
Each comparison of a prediction $\hat{y}$ to the corresponding ground truth $y$ yields either true positive $tp=\hat{y}\land y$, true negative $tn=\neg\hat{y}\land\neg y$, false positive $fp=\hat{y}\land\neg y$, or false negative $fn=\neg\hat{y}\land y$. Using these, binary classification metrics are defined: accuracy $a=\frac{tp+tn}{tp+tn+fp+fn}$, precision $p=\frac{tp}{tp+fp}$, recall $r=\frac{tp}{tp+fn}$, F1-score $f_{1}=\frac{2tp}{2tp+fp+fn}$, and Intersection over Union (IoU) $iou=\frac{tp}{tp+fn+tp}$.
As the regions of interest are relatively small with most of the queries, true negative predictions dominate the predictions. Therefore, accuracy cannot be used as a metric, as it incorporates true negative predictions. The other metrics are not affected by true negative predictions, so we use them as the metrics for binary classification tasks in this work.
IV-A2 Instance-based queryability
The second method evaluates the capability of the map to detect and retrieve all matching objects for the query. Each instance is predicted to match or not match the query; therefore, this measures the coverage of matches within an object rather than over the whole map.
While some latent semantic maps propose a way to perform instance segmentation, not all do, so measuring instance segmentation performance is outside of the scope of this work. Instead, we assume given ground-truth instances; these can be obtained through many different approaches, using latent semantic maps alone or combining them with other maps or modalities.
The method consists of the following steps:
  1. i)
    Given $\mathcal{M}$, $\mathcal{Q}$, and ground truth instance segmentation $\mathcal{I}$, each map $m\in\mathcal{M}$ is queried with the each query $q\in\mathcal{Q}$, each query yielding binary predicted mask $\hat{\textbf{y}}_{q}$.
  2. ii)
    For each object instance $i\in\mathcal{I}$, the corresponding voxels $\hat{\textbf{i}}_{i,q}\subset\hat{\textbf{y}}_{q}$ are selected from the predicted map.
  3. iii)
    If the majority of the voxels in $\hat{\textbf{i}}_{i,q}$ are $true$, the prediction $\hat{y}_{i,q}$ is $true$, otherwise $false$. The predictions for each instance for each query combined form the set of predictions $\hat{Y}_{I}=\{\hat{y}_{i,q}\forall i\in\mathcal{I};q\in\mathcal{Q}\}$.
  4. iv)
    Similar to the previous method, the true map $m_{q}$ is created. If the instance $i$ on the true map answers the query $q$, the true label $y_{i,q}=true$, otherwise $false$. The true labels for each instance for each query form the set of all true labels denoted $Y_{I}$.
  5. v)
    The binary classification metrics are calculated between $Y_{I}$ and $\hat{Y}_{I}$.
From $Y_{I}$ and $\hat{Y}_{I}$, the same metrics presented for the voxel-based queryability are computed for each instance.
IV-B Evaluating consistency
We further subdivide the consistency of the map into two distinct aspects, intra-map consistency and inter-map consistency, and propose a method for evaluating each.
IV-B1 Intra-map consistency
Intra-map consistency captures the similarity of embeddings across the voxels within a map, sharing semantic meaning. The hypothesis is that embeddings sharing semantic meaning are clustered together in the latent space to allow the separability of different concepts. While each object instance is in some sense unique, a consistent set of embeddings represents an abstract base class to which the instances belong.
The method consists of the following steps:
  1. i)
    Given a map $m$, and semantic label $l$, $\mathcal{E}_{l}$ is the set of embeddings corresponding to voxels in the map $m$ where the ground truth semantic label is $l$, and $\mathcal{E}_{m}$ is the set of all embeddings in the map $m$. We form a set of tuples $\mathcal{T}=\{(\mathcal{E}_{l},\mathcal{E}_{m})\forall m\in\mathcal{M};l\in\mathcal{L}\}$, where $\mathcal{L}$ is a closed-set semantic label vocabulary. For computational efficiency and to no considerable change in results, in practice, the sets $\mathcal{E}_{l}$ and $\mathcal{E}_{m}$ are subsampled. We use a subsampling ratio of $0.1$ in this work.
  2. ii)
    The average absolute deviation $d_{l}^{m}$ is calculated for $\mathcal{E}_{l}$, and $d_{m}^{m}$ for $\mathcal{E}_{m}$, for each tuple $t\in\mathcal{T}$. The average absolute deviation $d_{a}$ measures the statistical dispersion of a set, and given a set of points =$\mathcal{E}=\{e_{1},...,e_{n}\}$ it is defined as
    $d_{a}=\frac{1}{n}\sum^{n}_{j=1}|f_{d}(e_{j},\bar{\mathcal{E}})|,$ (1)
    where $f_{d}$ is a deviation metric and $\bar{\mathcal{E}}$ a central point of set $\mathcal{E}$. We use the mean of the embeddings as the central point and cosine distance as the deviation metric, which is consistent with the use of cosine distance loss for CLIP.
  3. iii)
    The intra-map consistency ratio $c_{l}^{m}=\frac{d_{l}^{m}}{d_{m}^{m}}$ is computed for each tuple $t\in\mathcal{T}$, with label $l$ in map $m$. This ratio represents the distinguishability of the class compared to the map average.
IV-B2 Inter-map consistency
Inter-map consistency captures the similarity of voxels with the same semantic label across different maps. The hypothesis is that embeddings within the same label are closer to each other across maps than to those with different labels with respect to a distance metric. This would imply that the objects retain their distinctiveness across maps, and therefore, the system generalizes better across different environments.
The method consists of the following steps:
  1. i)
    Given the set of tuples $\mathcal{T}$, constructed according to Section IV-B1.
  2. ii)
    For all pairs of tuples $\big{(}(\mathcal{E}_{l,1},\mathcal{E}_{m,1}),$ $(\mathcal{E}_{l,2},\mathcal{E}_{m,2})\big{)}$, $\mathcal{E}_{m,1}\neq\mathcal{E}_{m,2}$, parametric Wasserstein 2-distance $d_{w}$ is calculated between $\mathcal{E}_{l,1}$ and $\mathcal{E}_{l,2}$.
The Wasserstein $p$-distance is computationally heavy for large sets of high-dimensional embeddings. For this reason, given an n-dimensional embedding $e\in\mathbb{R}^{n}$, we approximate the set of embeddings $\mathcal{E}$ with an $n$-dimensional normal distribution $\mathcal{N}(\mu,P)$. This allows us to have a closed-form solution for the Wasserstein 2-distance
$d_{w}(\mathcal{E}_{1},\mathcal{E}_{2})\approx\lVert\mu_{1}-\mu_{2}\rVert^{2}_{2}+\text{tr}\big{(}P_{1}+P_{2}-2(P_{2}^{\frac{1}{2}}P_{1}P_{2}^{\frac{1}{2}})^{\frac{1}{2}}\big{)}.$ (2)
TABLE I: Average results of the voxel- and instance-based queryability tests over 10 maps.
Voxel-based Instance-based
Method Encoder F1 Precision Recall IoU F1 Precision Recall IoU
OpenScene LSeg 0.623 0.617 0.630 0.457 0.258 0.252 0.264 0.149
OpenScene OpenSeg 0.580 0.575 0.587 0.412 0.261 0.255 0.267 0.150
VLMaps LSeg 0.498 0.401 0.661 0.332 0.257 0.205 0.347 0.147
VLMaps OpenSeg 0.393 0.287 0.625 0.246 0.202 0.144 0.342 0.112
V Experiments
The main motivation of the experiments is to demonstrate the proposed benchmark. The two main questions the benchmark aims to answer with the experiments are:
  1. 1.
    How queryable state-of-the-art latent semantic maps are?
  2. 2.
    How consistent are their visual-language embeddings within and across maps?
To answer these questions, we evaluated two state-of-the-art methods, VLMaps [8] and OpenScene [11], in a series of experiments.
V-A Data set
The data set used in the benchmark is the Matterport3D data set [12], which consists of houses captured with an RGBD camera and ground truth semantic segmentation. The data set was used in the original work of VLMaps, in which they selected a sample of 10 sequences from the data set. In the benchmark, we use the same sequences and the same set of 42 labels forming the set $\mathcal{L}$; their names are also used in the following experiments as the query set $\mathcal{Q}$ to leverage the semantic ground truth provided with the dataset [^2] Supplementary material including the list of labels will be available in arXiv and link provided here upon acceptance of this work..
Since Matterport3D does not provide ground truth instance segmentation, instances used as ground truth were segmented using a region growing algorithm based on the ground truth semantics [32]. Each voxel is initialized as a seed cluster; then, region growing steps are performed, where the labels of all neighboring clusters are compared. If the labels are the same, the clusters are joined. Otherwise, they are not. This step is iterated until no more clusters can be joined or a maximum iteration limit is reached.
V-B Parametrization of the methods
Instead of using the 2D mapping method proposed in the original work [8], the VLMaps map was created using the new 3D mapping method available at the repository of the original authors [33]. This choice was made to create a fair comparison using state-of-the-art methods. The central idea is the same, except the 3D map avoids the problems when aggregating the embeddings along the $z$-axis. All of the parameters were the default used by the original authors. The 3D grid voxel size used is $0.05$ m, as proposed by the authors.
The 3D mapping method was also extended to create semantic ground truth maps from the Matterport3D data. The VLMaps use Habitat simulator [34, 35, 36] for camera measurements. Similarly to the RGB images, we create semantic images and backproject them to the 3D environment.
When queried with a query $q$, VLMaps creates a list of queries by setting $q$ and string ”other” into a list of phrases such as ”A photo of …” and ”there is … in the scene”. Then, the most similar query is selected from the list for each voxel. The binary mask is $true$ for voxels where the most similar query is a phrase containing $q$; otherwise, it is $false$. The binary mask is then dilated to encompass whole object instances.
OpenScene maps were created using the parameters proposed in the original paper. We use the proposed 2D-3D ensemble features. The 3D grid voxel size used is $0.02$ m, as proposed by the authors. OpenScene does not provide a way to query open-vocabulary binary masks. Because we use the set of labels as the queries, the binary query mask is created by comparing the equality of the voxel labels and the query.
Both methods used the same pre-trained LSeg and OpenSeg encoder provided by the original authors. Both encoders are based on the CLIP backbone; LSeg uses the CLIP-ViT-B/32 backbone with 512-dimensional embeddings, whereas OpenSeg uses the CLIP-ViT-L/14 with 768-dimensional embeddings.
Figure 2: The intra-map consistency ratios of the methods as histograms, with a kernel density estimation, which is smoothed with a Gaussian kernel. The mean of the distribution is depicted with a vertical line of corresponding color.
((a)) OpenScene with LSeg
((b)) OpenScene with OpenSeg
((c)) VLMaps with LSeg
((d)) VLMaps with OpenSeg
Figure 3: The distribution of Wasserstein distances of matching and non-matching labels are shown in orange and blue, respectively. The median of each distribution is presented with a dashed line.
((a)) OpenScene LSeg
((b)) OpenScene OpenSeg
((c)) VLMaps LSeg
((d)) VLMaps OpenSeg
Figure 4: The Wasserstein distances of the matching labels are depicted as orange bars and non-matching labels as blue bars on a logarithmic scale. The ten most common and uncommon labels are marked with red and blue text, respectively. The labels are sorted per image on the separability of the distance distribution according to Krusal-Wallis statistic [37]. The outliers of the boxplots have been omitted for clarity. The OpenScene distances, in (a) and (b), are multiple orders of magnitude larger, as the embeddings are an order of magnitude larger.
V-C Results
V-C1 Queryability
The results of the queryability benchmark are presented in Table I, where the F1-score, precision, recall, and IoU of the method and encoder combinations are presented for the voxel- and instance-based tests. Overall, OpenScene performs better than VLMaps with both encoders, based on the higher F1-scores and IoUs. The higher recall shown by VLMaps with LSeg is likely the result of the post-processing of the queries, where the query results are dilated to encompass whole objects. The instance-based results mostly align with the voxel-based results, except for the slightly decreased performance in all metrics.
While OpenScene with LSeg is better in the voxel-based experiment, both encoders have virtually the same performance in the instance-based experiment when used within OpenScene. This aligns with the findings presented in [11], where OpenScene with LSeg encoder was found to have better IoU.
Therefore, these results show that both the mapping method and choice of encoder matter in the map creation process. They also indicate that the 3D structure of the environment, included by OpenScene, contains information that can be leveraged in addition to the purely image-based creation of embeddings.
All methods have better recall than precision, which are usually tradeoffs. Better recall might be beneficial for queryability, especially in a navigation context: having more object candidates to explore than missing potential objects is better. This is especially visible in VLMaps, where the recall is considerably higher than precision. OpenScene has a more balanced approach, reflected in the high F1-Score, which is a good metric for the overall performance of the methods. IoU also conforms to the results, increasing their reliability.
As the instance classification is based on the prediction of the majority label, many objects seem to match the query only with a region representing less than half of the object. In this regard, the query methodology and, subsequently, the experiment setup could be improved: instead of classifying the results with binary masks, the minimum cosine distance could be used. This could allow the detection of instances with a small but significant match to the query. This would be relevant, for instance, when the object is partially occluded or unobserved. For example, if a sofa is partially covered with a blanket, the image-based methods could segment it as a blanket, but with the visible small region, it could still be recognized as a sofa. In this regard, the 3D understanding could prove extremely valuable: The geometry of the sofa might not significantly change when it is covered, so it can still be recognized as a sofa. When using open vocabulary, the description could even cover both objects: a sofa covered with a blanket.
Additionally, we qualitatively demonstrate the capability of the benchmark to process open-vocabulary queries. We use the VLMaps’ binary masking method without prompt engineering or dilation as the comparison method. From the results presented in Figure 1, where the matches are shown in orange and non-matches in blue, it can be seen that OpenScene query results are considerably less noisy and encompass entire objects, indicating that the features are more distinctive. While LSeg yields the objects more accurately, the more abstract queries, such as ”striped” or ”relax”, produce no results. OpenScene seems to be more sensitive, finding solutions to abstract queries but yielding many false positive results. While this is in line with the findings of the other experimental results, it might hint towards OpenSeg having a broader vocabulary.
V-C2 Consistency
The intra-map consistency
The results are presented in Figure 2, in which the intra-map consistency ratios of the methods are illustrated with histograms. Additionally, the means of data are shown with dashed lines. Notably, the ordering of the methods according to their intra-map consistency is the same as in the queryability experiment presented in Section V-C1. OpenScene has better consistency between the labels with each encoder, which is indicated by the lower means of the distributions. This indicates that OpenScene can cluster the embeddings of labels better using the ensemble features. VLMaps with OpenSeg has a long tail that continues beyond the figure, with several outliers.
However, there are multiple classes with a ratio above one. This means that the embeddings in a class are more different from each other than the map on average. This means these classes are challenging to cluster and hard to distinguish. However, these classes have, on average, 2-4% the number of voxels compared to the average number of voxels in the whole data set. For example, the two most represented categories in the tail are ”misc” and ”objects”, both relatively non-descriptive generic categories, and therefore, this could be attributed to the limited number of classes in the experiment.
The inter-map consistency
In Figure 3, the Wasserstein distances between matching, i.e., the label is the same, and non-matching sets of embeddings across all maps are shown. The medians of the data are shown with dashed lines. The ratio between the medians of matching and non-matching labels suggests that OpenScene separates the labels better than VLMaps, and LSeg separates them better than OpenSeg. The ratio of medians of OpenScene with LSeg is 9.21, OpenScene with OpenSeg is 3.67, VLMaps with LSeg 1.65, and VLMaps with OpenSeg 1.24, which is once again the exact ordering observed in Sections V-C1 and V-C2.
While the shapes of distributions are similar, OpenScene has much larger distances on average, even with matching labels. This results from the fact that the distilled features of OpenSeg are of a larger magnitude than the image features of both VLMaps and OpenScene image features. For example, the average norm of all OpenScene with OpenSeg embeddings in the map of the first sequence is 13.68 times larger than the average norm of all VLMaps with OpenSeg embeddings in the same map, and the maximum being 57.97 times larger than the maximum VLMaps feature.
Furthermore, in Figure 4, the distances of matching and non-matching distances are shown per class with orange and blue bars, respectively. The same phenomena can be seen: LSeg distinguishes the classes better than OpenSeg and OpenScene better than VLMaps, which is indicated by the separation of the blue and orange box plots. Additionally, the ten most common and uncommon categories in the Matterport3D data set are marked with red and blue names, respectively. VLMaps performs well in separating common classes and worse with uncommon classes, especially using LSeg, whereas OpenScene performs better across the whole set of labels.
When considering all these results, the combination of OpenScene and LSeg produces maps that are both more queryable and consistent, making it the method that currently can be expected to generalize better and lead to more consistent performance in downstream robotic applications.
VI Conclusion
Mobile robot maps must be able to capture the semantics of the environment for the robots to perform complex tasks. If the semantic representation is constrained to a closed set of classes, the robot’s ability to understand the environment is constrained to the same set. Latent semantic maps present a way to overcome this limitation and enable open-vocabulary scene understanding by representing the semantics in a latent space with a notion of distance.
In this work, we propose quantitative metrics to evaluate the quality of state-of-the-art latent semantic maps. In our experiments, OpenScene mapping using 3D scene structure with LSeg semantic embedding performed the best. This indicates that the 3D structure of the environment is useful for inferring the latent semantics. While LSeg performed best in the semantic segmentation setting, the qualitative demonstration hints that OpenSeg might be more sensitive to rarer and more abstract queries when properly thresholded.
The ability of visual-language maps to capture latent semantics is, however, still limited. Currently, fully supervised semantic segmentation methods outperform all latent semantic maps. The relatively low F1-scores and IoUs, especially in the instance-based test, indicate that there is still work to be done to address the problems of thresholding and clustering the query results to segment object instances.
A data set providing ground truth open-vocabulary semantics is direly needed. This would allow extending the evaluation presented here to adequately address the breadth of vocabulary and capture the full promise of VLMs.

### v2

@@
-Do Visual-Language Maps Capture Latent Semantics?
+Do Visual-Language Grid Maps Capture Latent Semantics?
 Abstract
-Visual-language models (VLMs) have recently been introduced in robotic mapping by using the latent representations, i.e., embeddings, of the VLMs to represent the natural language semantics in the map.
-The main benefit is moving beyond a small set of human-created labels toward open-vocabulary scene understanding.
-While there is anecdotal evidence that maps built this way support downstream tasks, such as navigation, rigorous analysis of the quality of the maps using these embeddings is lacking.
-We investigate two critical properties of map quality: queryability and consistency.
+Visual-language models (VLMs) have recently been introduced in robotic mapping using the latent representations, i.e., embeddings, of the VLMs to represent semantics in the map.
+They allow moving from a limited set of human-created labels toward open-vocabulary scene understanding, which is very useful for robots when operating in complex real-world environments and interacting with humans.
+While there is anecdotal evidence that maps built this way support downstream tasks, such as navigation, rigorous analysis of the quality of the maps using these embeddings is missing.
+In this paper, we propose a way to analyze the quality of maps created using VLMs.
+We investigate two critical properties of map quality: queryability and distinctness.
 The evaluation of queryability addresses the ability to retrieve information from the embeddings.
-We investigate two aspects of consistency: intra-map consistency and inter-map consistency.
-Intra-map consistency captures the ability of the embeddings to represent abstract semantic classes, and inter-map consistency captures the generalization properties of the representation.
-In this paper, we propose a way to analyze the quality of maps created using VLMs, which forms an open-source benchmark to be used when proposing new open-vocabulary map representations.
-We demonstrate the benchmark by evaluating the maps created by two state-of-the-art methods, VLMaps and OpenScene, using two encoders, LSeg and OpenSeg, using real-world data from the Matterport3D data set.
-We find that OpenScene outperforms VLMaps with both encoders, and LSeg outperforms OpenSeg with both methods.
+We investigate intra-map distinctness to study the ability of the embeddings to represent abstract semantic classes and inter-map distinctness to evaluate the generalization properties of the representation.
+We propose metrics to evaluate these properties and evaluate two state-of-the-art mapping methods, VLMaps and OpenScene, using two encoders, LSeg and OpenSeg, using real-world data from the Matterport3D data set.
+Our findings show that while 3D features improve queryability, they are not scale invariant, whereas image-based embeddings generalize to multiple map resolutions.
+This allows the image-based methods to maintain smaller map sizes, which can be crucial for using these methods in real-world deployments.
+Furthermore, we show that the choice of the encoder has an effect on the results.
+The results imply that properly thresholding open-vocabulary queries is an open problem.
 I Introduction
-Mobile robots must understand the geometry and semantics of the environment to accomplish complex tasks.
-The ability of maps to capture semantics has increased with the development of computer vision systems, from detecting single objects to creating pixel-wise semantic segmentation of images, yielding dense semantic maps with labels in each map cell [1].
-However, most semantic segmentation methods are trained to segment the image into a small pre-selected set of categories.
-This imposes challenges in real-world settings when a robot operates in an environment that the categories do not fully describe [2, 3].
-Visual-Language Models (VLMs) are networks that jointly train a visual and language encoder to learn a mapping from text and image inputs into a common visual-language latent space.
-Modern, efficient transformer architectures [4], such as CLIP [5], are trained with large data sets, which enables these models to effectively have an open vocabulary, meaning visual semantics can be matched to natural language instead of a set of selected symbols.
-Therefore, the exploitation of VLMs in robotic mapping could allow richer semantics to be represented and to overcome the difficulty of adapting to new environments.
-While this work focuses on maps that encode semantics using VLMs, mapping methods have been proposed that can incorporate other modalities as well, such as audio [6].
-We will collectively refer to all these maps as latent semantic maps.
-Figure 1: Maps built from visual-language models have the ability to represent complex semantics, encompassing both the class of objects and their properties.
-We propose a benchmark for these type of maps and evaluate the quality of different state-of-the-art methods according to the consistency of their representation and their ability to be queried.
+Robots must understand the geometry and semantics of the environment to accomplish complex tasks.
 The semantic information stored in the map must be able to be queried.
 While complex queries have been long possible in semantic maps by inference using object ontologies [1], semantic labels support only binary comparison as they are enumerations with no notion of distance.
-The promise of latent semantic maps is that queries can be evaluated by directly comparing their embeddings to the embeddings of the map with a similarity metric, yielding virtually unrestricted query vocabulary.
+This imposes challenges in real-world settings when a robot operates in an environment that the categories do not fully describe [2, 3].
+Figure 1: Maps built from visual-language models have the ability to represent complex semantics, encompassing both the class of objects and their properties.
+We propose a way to evaluate the quality of different state-of-the-art methods according to the distinctness of their semantic concepts and their ability to be queried.
+Lately, grid map methods have been proposed that use Visual-Language Model (VLM) embeddings as semantics, which we will collectively refer to as latent semantic maps.
+Their promise is that queries can be evaluated by directly comparing the query embedding to the embeddings of the map with a similarity metric, yielding virtually unrestricted query vocabulary.
 Several successful algorithms have been built on VLMs that can perform mobile robot tasks, such as navigation [4, 5, 6].
 However, the analysis from the robotic mapping perspective is left anecdotal.
-In this work, we propose a way to evaluate the quality of latent semantic maps.
-We consider that three aspects of the map capture its quality: queryability, consistency, and the breadth of the vocabulary.
-We concentrate on evaluating queryability and consistency because, to the best of our knowledge, no data sets exist to evaluate robotic maps that adequately capture open vocabulary semantics.
-Furthermore, we evaluate the consistency of the representation, i.e., the embeddings themselves.
-Understanding the possibilities and limitations of the representation is essential for bringing the methods into real-world applications.
-Having comparable metrics for the quality of the latent semantic maps is vital for measuring the progress in this new research field, as this allows for evaluating new methods against the state-of-the-art.
-For this reason, we provide an open-source benchmark [^1] The source code will be made publicly available upon acceptance of this work.
-against which any new open vocabulary semantic representation can be tested.
-To demonstrate our benchmark, we evaluate two state-of-the-art methods: VLMaps [8] and OpenScene [11], using the Matterport3D [12] data set.
+This work aims to bridge this gap by analyzing the performance of two state-of-the-art grid mapping methods, VLMaps [4] and OpenScene [7].
+We propose two metrics for this analysis: queryability and distinctness of semantic concepts.
+Queryability measures the ability to find matches for queries from the representation, while distinctness measures the ability to distinguish different abstract object classes.
+Furthermore, we explore the practical considerations of the methods by comparing changes in queryability with varying resolutions, as the map size is a function of the resolution of the map.
 The main contributions of this paper are:
   1. i)
-    We define queryability and consistency of representation as measures of mapping quality for latent semantic maps and propose metrics to quantitatively evaluate these properties;
+    We define queryability and distinctness of representation as measures of mapping quality for latent semantic maps and propose metrics for quantitatively evaluating the distinctness;
   2. ii)
-    We provide an open-source benchmark software for latent semantic maps, enabling evaluation of the progress in this new field of research;
+    We provide a comprehensive analysis of two state-of-the-art methods using the proposed metrics;
   3. iii)
-    We provide a comprehensive analysis of two state-of-the-art methods using the proposed benchmark.
+    We evaluate the behavior of the maps with different resolutions;
+  4. iv)
+    We make our evaluation software publicly available [^1] The source code will be made publicly available upon acceptance of this work., enabling evaluation of the progress in this new field of research.
 II Related work
-II-A Maps created with visual-language embeddings
-Creating maps using embeddings of VLMs, especially CLIP [5], has recently gained attention, with the proposal of methods such as VLMaps [8], OpenScene [11], NLMaps [13], and Uni-Fusion [14].
-VLMaps and OpenScene create a dense grid map, where each cell is associated with a single VLM embedding.
+II-A Grid maps created with visual-language embeddings
+Creating maps using embeddings of VLMs, especially CLIP [8], has recently gained attention, with the proposal of methods such as VLMaps [4], OpenScene [7], NLMaps [9], Uni-Fusion [10], ConceptFusion [11], ConceptGraphs [12], and HOV-SG [13].
+The most common type of map used in robotics is a dense grid map.
+VLMaps and OpenScene have augmented these maps with open-vocabulary semantics by associating each cell with a single VLM embedding.
 The embeddings are created from RGBD images with a language-driven semantic segmentation encoder, such as LSeg [14] or OpenSeg [15], both based on CLIP, and back-projected to the 3D map.
 Each cell is represented as the mean of the embeddings projected to the cell.
-In the original work, the 3D map in VLMaps is projected to 2D by averaging the embeddings in the $z$-axis.
+In the original work, the 3D map in VLMaps is projected to 2D by averaging the embeddings on the $z$-axis.
 Additionally, OpenScene learns an encoder that directly produces CLIP embeddings from the 3D point cloud.
-Combining these two approaches, they propose joint 2D-3D ”ensemble” features as the final representations in the map.
-Unlike VLMaps and OpenScene, NLMaps creates a feature-based map, where each feature is associated with an VLM embedding.
+Combining these two approaches, OpenScene proposes joint 2D-3D “ensemble” features as the final representations in the map.
+Other methods have been proposed, which do not use grid maps or do not produce embeddings on each grid cell.
+NLMaps creates a feature-based map, where each feature is associated with an VLM embedding.
 Uni-Fusion proposes a general method to create continuous maps from any perceptual data modality, including VLM embeddings.
-In this paper, we focus on VLMaps and OpenScene because the NLMaps and Uni-Fusion source code were not made fully publicly available at the time of writing.
+ConceptFusion builds a dense point cloud map with an embedding for each point, whereas ConceptGraphs and HOV-SG have dense point cloud maps with one embedding per object.
+The extension of the evaluation of representations with object-level embeddings is left for future works.
 II-B Open-vocabulary queries
@@
 In contrast, in this work, we evaluate queryability such that the evaluation is not restricted to semantic segmentation of the map, which imposes that the map must be partitioned.
-Instead, each query is evaluated separately.
-II-C Analysis of maps
-When assessing the quality of robotic maps, most existing works concentrate on evaluating geometry using methods ranging from simple feature descriptors [23], frequency-based methods [24], to probabilistic measures [25].
-Evaluating the geometric and semantic properties jointly is only relevant when the geometry and semantics are jointly estimated, e.g., in semantic Simulatenous Localization and Mapping (SLAM).
+Instead, each query is evaluated separately, as it would be in real applications.
+II-C Analysis of quality of maps
+Most existing works assessing the quality of robotic maps concentrate on evaluating geometry using methods ranging from simple feature descriptors [22], frequency-based methods [23], to probabilistic measures [24].
+Evaluating the geometric and semantic properties jointly is only relevant when the geometry and semantics are jointly estimated, e.g., in semantic Simultaneous Localization and Mapping (SLAM).
 Often, the metrics are qualitative, but quantitative metrics such as the localization accuracy or map reconstruction quality compared to ground truth objects are used as well [25, 26, 27].
@@
 III Problem statement
-We aim to create a benchmark to evaluate the quality of latent semantic maps.
-We only consider methods where the estimation of geometry and semantics are disjoint, so we concentrate on evaluating the semantic quality of the maps.
-We believe that three aspects of the map capture its semantic quality: queryability, consistency, and the breadth of the vocabulary.
-We concentrate on evaluating queryability and consistency because, to the best of our knowledge, no data sets exist to evaluate robotic maps that adequately capture open-vocabulary semantics.
-The benchmark must evaluate queryability, as it is the primary way to retrieve information from the representation and, therefore, acts as the measure of two things: first, indirectly, the ability of the representation to contain relevant information, and second, the accessibility of that information.
-It also must address consistency, as consistent performance within and between measuring runs, times, environments, and sensors is desirable.
+In this study, we evaluate the quality of latent semantic maps.
+We only consider methods where the estimation of geometry and semantics are disjoint, so we concentrate on evaluating the semantic quality of the maps, which we consider to be captured by two properties: queryability and distinctness.
+Given a map $m_{i}$, an open-vocabulary query $q$, and a query function $\varphi$, a binary mask $b^{i}_{q}=\varphi(m_{i},q)$ is acquired.
+Given the ground truth binary mask $b^{i}_{t}$ for the same query, the queryability $Q$ with respect to $q$ and a metric $F_{q}$ is defined as $Q^{i}_{q}=F_{q}(b^{i}_{q},b^{i}_{t})$.
+Given a set $\mathcal{S}^{i}_{l}$ of semantic representations of a class $l$ in map $m_{i}$, distinctness $D$ between two labels $l$ and $l^{\prime}$, from maps $m_{i}$ and $m_{i^{\prime}}$, respectively, and with respect to a metric $F_{d}$, is defined as $D^{i,\,l}_{i^{\prime},\,l^{\prime}}=F_{d}(\mathcal{S}^{i}_{l},\mathcal{S}^{i^{\prime}}_{l^{\prime}})$.
 IV Methods
 IV-A Evaluating queryability
-Because the environment consists both of background regions (e.g., floor and ceiling) and discrete object instances (e.g., a chair or a table), we propose evaluating queryability in two ways: by measuring the voxel-based coverage of the query results and the instance retrieval capability.
-IV-A1 Voxel-based queryability
-The first method evaluates the overall matching between the query results compared to the ground truth in a binary classification setting.
+Queryability captures the ability to retrieve information from the embeddings, i.e., to find the parts of the map that answer the query.
+The hypothesis is that from a map with high queryability, it is possible to find accurate matches to a wide range of queries.
+Queryability is evaluated by comparing the query results and ground truth in a binary classification setting.
 This property cannot be evaluated using standard multi-class classification, as queries do not form a partition of the map.
-Each query produces a query result, a binary mask, over the whole map, and the same voxel might be a match of multiple queries.
-This is evident from the open-vocabulary perspective; e.g., a sofa might match queries ”sofa”, ”couch”, ”place to sit”, and ”soft”.
+Each query result is a binary mask over the whole map, and the same voxel might be a match of multiple queries.
+This is evident from the open-vocabulary perspective; e.g., a sofa might match queries “sofa”, “couch”, “place to sit”, and “soft”.
 The method consists of the following steps:
@@
 The query result of a single query is a binary segmentation mask of the map, $\hat{\textbf{y}}_{q}=\{\hat{y}_{1},\dots,\hat{y}_{N}\}$, consisting of $N$ voxel masks $\hat{y}\in\{true,false\}$.
-Combined, the query results form the set of voxel-based predictions $\hat{Y}_{V}=\{\hat{\textbf{y}}_{q}\forall q\in\mathcal{Q}\}$.
+Combined, the query results form the set of voxel-based predictions $\hat{Y}_{V}=\{\hat{\textbf{y}}_{q}|q\in\mathcal{Q}\}$.
   3. iii)
     For each query, the true mask $\textbf{y}_{q}$ is created from the ground truth map $m_{g}$ such, that the voxels answering query $q$ are $true$, others $false$.
-The true masks form a set of true labels $Y_{V}=\{\textbf{y}_{q}\forall q\in\mathcal{Q}\}$.
+The true masks form a set of true labels $Y_{V}=\{\textbf{y}_{q}|q\in\mathcal{Q}\}$.
   4. iv)
     The binary classification metrics are calculated between $Y_{V}$ and $\hat{Y}_{V}$.
-Each comparison of a prediction $\hat{y}$ to the corresponding ground truth $y$ yields either true positive $tp=\hat{y}\land y$, true negative $tn=\neg\hat{y}\land\neg y$, false positive $fp=\hat{y}\land\neg y$, or false negative $fn=\neg\hat{y}\land y$.
-Using these, binary classification metrics are defined: accuracy $a=\frac{tp+tn}{tp+tn+fp+fn}$, precision $p=\frac{tp}{tp+fp}$, recall $r=\frac{tp}{tp+fn}$, F1-score $f_{1}=\frac{2tp}{2tp+fp+fn}$, and Intersection over Union (IoU) $iou=\frac{tp}{tp+fn+tp}$.
-As the regions of interest are relatively small with most of the queries, true negative predictions dominate the predictions.
+We use precision, recall, f1-score, and Intersection over Union (IoU) as metrics.
+We acquire the queryability by averaging them over all of the maps and queries with respect to each metric $Q^{\mathcal{M}}_{\mathcal{Q}}=F_{q}(Y_{V},\hat{Y}_{V}),\>F_{q}\in\{p,r,f_{1},iou\}$.
+With most queries, the regions of interest are relatively small, so true negatives dominate the predictions.
 Therefore, accuracy cannot be used as a metric, as it incorporates true negative predictions.
-The other metrics are not affected by true negative predictions, so we use them as the metrics for binary classification tasks in this work.
-IV-A2 Instance-based queryability
-The second method evaluates the capability of the map to detect and retrieve all matching objects for the query.
-Each instance is predicted to match or not match the query; therefore, this measures the coverage of matches within an object rather than over the whole map.
-While some latent semantic maps propose a way to perform instance segmentation, not all do, so measuring instance segmentation performance is outside of the scope of this work.
-Instead, we assume given ground-truth instances; these can be obtained through many different approaches, using latent semantic maps alone or combining them with other maps or modalities.
+The selected metrics are not affected by this problem.
+IV-B Evaluating distinctness
+We further subdivide the distinctness of the map into two distinct aspects, intra-map distinctness, and inter-map distinctness, and propose a method for evaluating each.
+IV-B1 Intra-map distinctness
+Intra-map distinctness captures the similarity of embeddings across the voxels within a map, sharing semantic meaning.
+The hypothesis is that embeddings sharing semantic meaning are clustered together in the latent space, allowing the distinguishing of different concepts.
 The method consists of the following steps:
   1. i)
-    Given $\mathcal{M}$, $\mathcal{Q}$, and ground truth instance segmentation $\mathcal{I}$, each map $m\in\mathcal{M}$ is queried with the each query $q\in\mathcal{Q}$, each query yielding binary predicted mask $\hat{\textbf{y}}_{q}$.
-  2. ii)
-    For each object instance $i\in\mathcal{I}$, the corresponding voxels $\hat{\textbf{i}}_{i,q}\subset\hat{\textbf{y}}_{q}$ are selected from the predicted map.
-  3. iii)
-    If the majority of the voxels in $\hat{\textbf{i}}_{i,q}$ are $true$, the prediction $\hat{y}_{i,q}$ is $true$, otherwise $false$.
-The predictions for each instance for each query combined form the set of predictions $\hat{Y}_{I}=\{\hat{y}_{i,q}\forall i\in\mathcal{I};q\in\mathcal{Q}\}$.
-  4. iv)
-    Similar to the previous method, the true map $m_{q}$ is created.
-If the instance $i$ on the true map answers the query $q$, the true label $y_{i,q}=true$, otherwise $false$.
-The true labels for each instance for each query form the set of all true labels denoted $Y_{I}$.
-  5. v)
-    The binary classification metrics are calculated between $Y_{I}$ and $\hat{Y}_{I}$.
-From $Y_{I}$ and $\hat{Y}_{I}$, the same metrics presented for the voxel-based queryability are computed for each instance.
-IV-B Evaluating consistency
-We further subdivide the consistency of the map into two distinct aspects, intra-map consistency and inter-map consistency, and propose a method for evaluating each.
-IV-B1 Intra-map consistency
-Intra-map consistency captures the similarity of embeddings across the voxels within a map, sharing semantic meaning.
-The hypothesis is that embeddings sharing semantic meaning are clustered together in the latent space to allow the separability of different concepts.
-While each object instance is in some sense unique, a consistent set of embeddings represents an abstract base class to which the instances belong.
-The method consists of the following steps:
-  1. i)
-    Given a map $m$, and semantic label $l$, $\mathcal{E}_{l}$ is the set of embeddings corresponding to voxels in the map $m$ where the ground truth semantic label is $l$, and $\mathcal{E}_{m}$ is the set of all embeddings in the map $m$.
-We form a set of tuples $\mathcal{T}=\{(\mathcal{E}_{l},\mathcal{E}_{m})\forall m\in\mathcal{M};l\in\mathcal{L}\}$, where $\mathcal{L}$ is a closed-set semantic label vocabulary.
-For computational efficiency and to no considerable change in results, in practice, the sets $\mathcal{E}_{l}$ and $\mathcal{E}_{m}$ are subsampled.
+    Given the $i$-th map $m_{i}$, and semantic label $l$, $\mathcal{E}_{l}$ is the set of embeddings corresponding to voxels in the map $m_{i}$ where the ground truth semantic label is $l$, and $\mathcal{E}_{i}$ is the set of all embeddings in the map $m_{i}$.
+We form a set of tuples $\mathcal{T}=\{(\mathcal{E}_{l},\mathcal{E}_{i})|m_{i}\in\mathcal{M},l\in\mathcal{L}\}$, where $\mathcal{L}$ is a closed-set semantic label vocabulary.
+For computational efficiency and to no considerable change in results, in practice, the sets $\mathcal{E}_{l}$ and $\mathcal{E}_{i}$ are subsampled.
+We subsample each label separately to preserve the label distribution.
 We use a subsampling ratio of $0.1$ in this work.
   2. ii)
-    The average absolute deviation $d_{l}^{m}$ is calculated for $\mathcal{E}_{l}$, and $d_{m}^{m}$ for $\mathcal{E}_{m}$, for each tuple $t\in\mathcal{T}$.
+    The average absolute deviation $d_{l}^{i}$ is calculated for $\mathcal{E}_{l}$, and $d_{\mu}^{i}$ for $\mathcal{E}_{i}$, for each tuple $t\in\mathcal{T}$.
 The average absolute deviation $d_{a}$ measures the statistical dispersion of a set, and given a set of points =$\mathcal{E}=\{e_{1},...,e_{n}\}$ it is defined as
@@
   3. iii)
-    The intra-map consistency ratio $c_{l}^{m}=\frac{d_{l}^{m}}{d_{m}^{m}}$ is computed for each tuple $t\in\mathcal{T}$, with label $l$ in map $m$.
-This ratio represents the distinguishability of the class compared to the map average.
-IV-B2 Inter-map consistency
-Inter-map consistency captures the similarity of voxels with the same semantic label across different maps.
-The hypothesis is that embeddings within the same label are closer to each other across maps than to those with different labels with respect to a distance metric.
-This would imply that the objects retain their distinctiveness across maps, and therefore, the system generalizes better across different environments.
+    The intra-map distinctness ratio $c_{l}^{i}=\frac{d_{l}^{i}}{d_{\mu}^{i}}$ is computed for each tuple $t\in\mathcal{T}$, with label $l$ in map $m_{i}$.
+This ratio represents the distinctness $D^{i,\,l}_{i,\,\mathcal{L}}$ of the class $l$ compared to the map average.
+IV-B2 Inter-map distinctness
+Inter-map distinctness captures the similarity of voxels with the same semantic label across different maps.
+The hypothesis is that embeddings within the same label are closer to each other across maps than to those with different labels with respect to a distance metric, i.e., $D^{i,\,l}_{j,\,l}<D^{i,\,l}_{i,\,l^{\prime}}$.
+This would imply that the object classes retain their distinctness across maps, and therefore, the system generalizes better across different environments.
 The method consists of the following steps:
@@
   2. ii)
-    For all pairs of tuples $\big{(}(\mathcal{E}_{l,1},\mathcal{E}_{m,1}),$ $(\mathcal{E}_{l,2},\mathcal{E}_{m,2})\big{)}$, $\mathcal{E}_{m,1}\neq\mathcal{E}_{m,2}$, parametric Wasserstein 2-distance $d_{w}$ is calculated between $\mathcal{E}_{l,1}$ and $\mathcal{E}_{l,2}$.
-The Wasserstein $p$-distance is computationally heavy for large sets of high-dimensional embeddings.
+    For all pairs of tuples $\big{(}(\mathcal{E}_{l_{1}},\mathcal{E}_{i_{1}}),$ $(\mathcal{E}_{l_{2}},\mathcal{E}_{i_{2}})\big{)}$, ${i_{1}}\neq{i_{2}}$, parametric Wasserstein 2-distance $d_{w}$ is calculated between $\mathcal{E}_{l_{1}}$ and $\mathcal{E}_{l_{2}}$, representing distinctness $D^{i_{1},\,l_{1}}_{i_{2},\,l_{2}}$.
+We subsample $\mathcal{E}_{l_{1}}$ and $\mathcal{E}_{l_{2}}$ similarly to Section IV-B1.
+If the embeddings are normalized, they can be thought of as high-dimensional probability distributions.
+To measure the distance between the distributions of $\mathcal{E}_{l,1}$ and $\mathcal{E}_{l,2}$, we selected the Wasserstein $p$-distance, which is a common metric of the distance between probability distributions.
+However, it is computationally heavy for large sets of high-dimensional embeddings.
 For this reason, given an n-dimensional embedding $e\in\mathbb{R}^{n}$, we approximate the set of embeddings $\mathcal{E}$ with an $n$-dimensional normal distribution $\mathcal{N}(\mu,P)$.
@@
 $d_{w}(\mathcal{E}_{1},\mathcal{E}_{2})\approx\lVert\mu_{1}-\mu_{2}\rVert^{2}_{2}+\text{tr}\big{(}P_{1}+P_{2}-2(P_{2}^{\frac{1}{2}}P_{1}P_{2}^{\frac{1}{2}})^{\frac{1}{2}}\big{)}.$ (2)
-TABLE I: Average results of the voxel- and instance-based queryability tests over 10 maps.
-Voxel-based Instance-based
+Figure 2: The mean map size in gigabits (Gb) as a function of the map cell size.
+TABLE I: Average results of the queryability tests over 10 maps.
+Semantic segmentation VLMaps querying
 Method Encoder F1 Precision Recall IoU F1 Precision Recall IoU
-OpenScene LSeg 0.623 0.617 0.630 0.457 0.258 0.252 0.264 0.149
-OpenScene OpenSeg 0.580 0.575 0.587 0.412 0.261 0.255 0.267 0.150
-VLMaps LSeg 0.498 0.401 0.661 0.332 0.257 0.205 0.347 0.147
-VLMaps OpenSeg 0.393 0.287 0.625 0.246 0.202 0.144 0.342 0.112
+OpenScene LSeg 0.645 0.639 0.651 0.480 0.174 0.096 0.897 0.095
+OpenScene OpenSeg 0.603 0.598 0.609 0.435 0.111 0.059 0.975 0.059
+VLMaps LSeg 0.541 0.529 0.553 0.375 0.425 0.297 0.757 0.271
+VLMaps OpenSeg 0.495 0.526 0.469 0.333 0.367 0.256 0.700 0.226
 V Experiments
-The main motivation of the experiments is to demonstrate the proposed benchmark. The two main questions the benchmark aims to answer with the experiments are:
+The three main questions that we aim to answer with the experiments are:
   1. 1.
@@
   2. 2.
-    How consistent are their visual-language embeddings within and across maps?
+    How distinct are their visual-language embeddings within and across maps?
+  3. 3.
+    How does the grid map resolution affect the queryability?
 To answer these questions, we evaluated two state-of-the-art methods, VLMaps [4] and OpenScene [7], in a series of experiments.
 V-A Data set
-The data set used in the benchmark is the Matterport3D data set [12], which consists of houses captured with an RGBD camera and ground truth semantic segmentation.
+The data set used in the study is the Matterport3D data set [31], which consists of houses captured with an RGBD camera and ground truth semantic segmentation.
 The data set was used in the original work of VLMaps, in which they selected a sample of 10 sequences from the data set.
-In the benchmark, we use the same sequences and the same set of 42 labels forming the set $\mathcal{L}$; their names are also used in the following experiments as the query set $\mathcal{Q}$ to leverage the semantic ground truth provided with the dataset [^2] Supplementary material including the list of labels will be available in arXiv and link provided here upon acceptance of this work..
+In this work, we use the same sequences and the same set of 40 labels [31] forming the set $\mathcal{L}$; their names are also used in the following experiments as the query set $\mathcal{Q}$ to leverage the semantic ground truth provided with the dataset [^2] Supplementary material including the list of labels will be available in arXiv and link provided here upon acceptance of this work..
 Since Matterport3D does not provide ground truth instance segmentation, instances used as ground truth were segmented using a region growing algorithm based on the ground truth semantics [32].
@@
 All of the parameters were the default used by the original authors.
-The 3D grid voxel size used is $0.05$ m, as proposed by the authors.
 The 3D mapping method was also extended to create semantic ground truth maps from the Matterport3D data.
@@
 Similarly to the RGB images, we create semantic images and backproject them to the 3D environment.
-When queried with a query $q$, VLMaps creates a list of queries by setting $q$ and string ”other” into a list of phrases such as ”A photo of …” and ”there is … in the scene”.
-Then, the most similar query is selected from the list for each voxel.
-The binary mask is $true$ for voxels where the most similar query is a phrase containing $q$; otherwise, it is $false$.
-The binary mask is then dilated to encompass whole object instances.
 OpenScene maps were created using the parameters proposed in the original paper.
-We use the proposed 2D-3D ensemble features.
-The 3D grid voxel size used is $0.02$ m, as proposed by the authors.
-OpenScene does not provide a way to query open-vocabulary binary masks.
-Because we use the set of labels as the queries, the binary query mask is created by comparing the equality of the voxel labels and the query.
+The main analysis uses the proposed 2D-3D ensemble features, but both 2D image features (“fusion”) and 3D point cloud features (“distill”) were created for the ablation study.
 Both methods used the same pre-trained LSeg and OpenSeg encoder provided by the original authors.
 Both encoders are based on the CLIP backbone; LSeg uses the CLIP-ViT-B/32 backbone with $512$-dimensional embeddings, whereas OpenSeg uses the CLIP-ViT-L/14 with $768$-dimensional embeddings.
-Figure 2: The intra-map consistency ratios of the methods as histograms, with a kernel density estimation, which is smoothed with a Gaussian kernel.
+V-C Querying the maps
+VLMaps proposes a method for performing open-vocabulary queries, which we will refer to as VLMaps querying.
+When the map is queried with a query $q$, the embeddings on the map are compared using cosine similarity to two embeddings, embedding of $q$ and embedding of string “other”.
+The binary classification mask is $true$ for voxels where the most similar query is $q$; otherwise, it is $false$.
+The binary mask is then post-processed by a sequence of binary closing, gaussian blurring, thresholding, and binary dilation to encompass whole object instances.
+OpenScene does not propose a method to perform open-vocabulary queries.
+In the original work, the map is evaluated in a semantic segmentation setting, and open-vocabulary queries are evaluated qualitatively.
+In our experiment, we implement the VLMaps querying method on OpenScene and use the same parametrization for both methods.
+Additionally, we evaluate both methods in a semantic segmentation setting, where a binary query mask is acquired by comparing the equality of the voxel labels on the map and the query label.
+Finally, VLMaps proposed to use prompt engineering by creating a list of queries by setting $q$ and string “other” into a list of phrases such as “A photo of …” and “there is … in the scene”.
+Then, the augmented lists of prompts are embedded, and the embeddings are averaged to acquire two embeddings, one for $q$, one for “other”, for comparison.
+We evaluate the effect of prompt engineering on queryability on VLMaps.
+V-D Resolution analysis
+For real-world applications, where the robot hardware capabilities might be limited, the memory footprint of the map can be crucial.
+Small cell sizes can lead to prohibitively large maps, and the size of a grid map is a function of the resolution of the map, as shown in Figure 2.
+OpenScene has smaller map sizes due to downsampling the images.
+To investigate the effects of cell size on the map quality, all of the experiments were performed with a set of resolutions $r\in[0.02,0.05,0.1,0.2]$ m.
+The resolutions proposed in the original works are $0.02$m and $0.05$m in OpenScene and VLMaps, respectively.
+If not explicitly stated otherwise, we show the results in a resolution of $0.02$m for fair comparison.
+VI Results
+VI-A Queryability
+The results of the queryability are presented in Table I, where the F1-score, precision, recall, and IoU of all of the method and encoder combinations are presented for the semantic segmentation and VLMaps querying.
+VI-A1 Semantic segmentation
+In semantic segmentation, overall, OpenScene performs better than VLMaps with both encoders, based on the higher F1-scores and IoUs.
+Furthermore, LSeg performs better with both methods than OpenSeg.
+Therefore, these results show that both the mapping method and choice of encoder matter in the map creation process.
+They also indicate that the 3D structure of the environment, included by OpenScene, contains information that can be leveraged in addition to the purely image-based creation of embeddings.
+Figure 3: The F1-score of the methods in semantic segmentation task as a function of the map cell size.
+VI-A2 VLMaps querying
+In the queryability experiment using VLMaps querying, both VLMaps methods performed substantially better than OpenScene.
+While the results between semantic segmentation and VLMaps querying are not directly comparable, overall low F1-scores indicate that comparison with a single label, “other”, is not sufficient.
+It is unlikely that any one label would suffice as a universal complement for arbitrary open-vocabulary queries to threshold the results accurately.
+In a semantic segmentation setting, when the label set is suitable to the data set, the larger set of comparison labels could alleviate this problem by allowing more points of comparison to threshold the query results.
+VI-A3 Prompt engineering
+One approach to improve the matching of text queries proposed in OpenScene and VLMaps is to use prompt engineering.
+In our experiment, prompt engineering improves the precision of VLMaps methods, yielding an average $0.059$ and $0.008$ increase in the F1-score on LSeg and OpenSeg, respectively.
+While OpenScene proposes to use a single prompt augmentation “a XX in a scene”, the set of prompt engineering phrases used in VLMaps did not work with OpenScene.
+This indicates that the encoders are sensitive to prompt engineering and that a single set of phrases might not work with all encoders.
+Therefore, while prompt engineering could improve the performance, it does not solve the problem of open-vocabulary query thresholding.
+VI-A4 Resolution
+The effect of map voxel size, i.e., the resolution, to the F1-score is shown in Figure 3.
+We present only the F1-score for presentation clarity, as the precision/recall ratio remains similar over the resolutions.
+The effect of resolution change is minimal in VLMaps as image features are naturally more scale-invariant, as objects of the same category are perceived from different distances in the training data.
+The 3D features used by OpenScene are not scale invariant, as the point cloud measurements are on an absolute scale.
+This leads the performance to decrease rapidly.
+As the difference between the training resolution and the map resolution grows, the number of out-of-distribution samples grows.
+This effect is especially pronounced in OpenScene using LSeg.
+VI-A5 Open-vocabulary
+Additionally, we qualitatively demonstrate our evaluation metrics with open-vocabulary queries.
+The queries, which are only used in this qualitative demonstration, are: “sofa”, “couch”, “white sofa”, “striped”, and “relax”.
+We use the VLMaps’ binary masking method without prompt engineering or dilation as the comparison method.
+From the results presented in Figure 1, where the matches are shown in orange and non-matches in blue, it can be seen that OpenScene query results are considerably less noisy and encompass entire objects, indicating that the features are more distinctive.
+While LSeg yields the objects more accurately, the more abstract queries, such as “striped” or “relax”, produce no results.
+OpenSeg seems to be more sensitive, finding solutions to abstract queries but yielding many false positive results.
+While this is in line with the findings of the other experimental results, it might hint towards OpenSeg having a broader vocabulary.
+Figure 4: The intra-map distinctness ratios of the methods as histograms, with a kernel density estimation, which is smoothed with a Gaussian kernel.
 The mean of the distribution is depicted with a vertical line of corresponding color.
@@
 The OpenScene distances, in (a) and (b), are multiple orders of magnitude larger, as the embeddings are an order of magnitude larger.
-V-C Results
-V-C1 Queryability
-The results of the queryability benchmark are presented in Table I, where the F1-score, precision, recall, and IoU of the method and encoder combinations are presented for the voxel- and instance-based tests.
-Overall, OpenScene performs better than VLMaps with both encoders, based on the higher F1-scores and IoUs.
-The higher recall shown by VLMaps with LSeg is likely the result of the post-processing of the queries, where the query results are dilated to encompass whole objects.
-The instance-based results mostly align with the voxel-based results, except for the slightly decreased performance in all metrics.
-While OpenScene with LSeg is better in the voxel-based experiment, both encoders have virtually the same performance in the instance-based experiment when used within OpenScene.
-This aligns with the findings presented in [11], where OpenScene with LSeg encoder was found to have better IoU.
-Therefore, these results show that both the mapping method and choice of encoder matter in the map creation process.
-They also indicate that the 3D structure of the environment, included by OpenScene, contains information that can be leveraged in addition to the purely image-based creation of embeddings.
-All methods have better recall than precision, which are usually tradeoffs.
-Better recall might be beneficial for queryability, especially in a navigation context: having more object candidates to explore than missing potential objects is better.
-This is especially visible in VLMaps, where the recall is considerably higher than precision.
-OpenScene has a more balanced approach, reflected in the high F1-Score, which is a good metric for the overall performance of the methods.
-IoU also conforms to the results, increasing their reliability.
-As the instance classification is based on the prediction of the majority label, many objects seem to match the query only with a region representing less than half of the object.
-In this regard, the query methodology and, subsequently, the experiment setup could be improved: instead of classifying the results with binary masks, the minimum cosine distance could be used.
-This could allow the detection of instances with a small but significant match to the query.
-This would be relevant, for instance, when the object is partially occluded or unobserved.
-For example, if a sofa is partially covered with a blanket, the image-based methods could segment it as a blanket, but with the visible small region, it could still be recognized as a sofa.
-In this regard, the 3D understanding could prove extremely valuable: The geometry of the sofa might not significantly change when it is covered, so it can still be recognized as a sofa.
-When using open vocabulary, the description could even cover both objects: a sofa covered with a blanket.
-Additionally, we qualitatively demonstrate the capability of the benchmark to process open-vocabulary queries.
-We use the VLMaps’ binary masking method without prompt engineering or dilation as the comparison method.
-From the results presented in Figure 1, where the matches are shown in orange and non-matches in blue, it can be seen that OpenScene query results are considerably less noisy and encompass entire objects, indicating that the features are more distinctive.
-While LSeg yields the objects more accurately, the more abstract queries, such as ”striped” or ”relax”, produce no results.
-OpenScene seems to be more sensitive, finding solutions to abstract queries but yielding many false positive results.
-While this is in line with the findings of the other experimental results, it might hint towards OpenSeg having a broader vocabulary.
-V-C2 Consistency
-The intra-map consistency
-The results are presented in Figure 2, in which the intra-map consistency ratios of the methods are illustrated with histograms.
+VI-B Distinctness
+VI-B1 Intra-map distinctness
+The results are presented in Figure 4, in which the intra-map distinctness ratios of the methods are illustrated with histograms.
 Additionally, the means of data are shown with dashed lines.
-Notably, the ordering of the methods according to their intra-map consistency is the same as in the queryability experiment presented in Section V-C1.
-OpenScene has better consistency between the labels with each encoder, which is indicated by the lower means of the distributions.
-This indicates that OpenScene can cluster the embeddings of labels better using the ensemble features.
-VLMaps with OpenSeg has a long tail that continues beyond the figure, with several outliers.
-However, there are multiple classes with a ratio above one.
-This means that the embeddings in a class are more different from each other than the map on average.
-This means these classes are challenging to cluster and hard to distinguish.
-However, these classes have, on average, 2-4% the number of voxels compared to the average number of voxels in the whole data set.
-For example, the two most represented categories in the tail are ”misc” and ”objects”, both relatively non-descriptive generic categories, and therefore, this could be attributed to the limited number of classes in the experiment.
-The inter-map consistency
-In Figure 3, the Wasserstein distances between matching, i.e., the label is the same, and non-matching sets of embeddings across all maps are shown.
+LSeg has better distinctness between the labels with each method, which is indicated by the lower means of the distributions.
+This indicates that LSeg features of semantic classes are more distinct from each other than in OpenSeg.
+Furthermore, the 3D ensemble features of OpenScene help to make the features slightly more distinct, as OpenScene has lower means than VLMaps with comparable encoders.
+More distinct features, with lower intra-map distinctness ratios, should be preferable, as that indicates better queryability performance.
+The intuition is that in a best-case scenario, where the query is very accurate, i.e., close to the desired map features in the embedding space, tightly clustered classes are retrieved as a whole, which should lead to better results.
+More spread-out embeddings are more likely to return partial and sporadic results.
+There are multiple classes with a ratio above one, which means that the embeddings in those classes are more different from each other than the map on average.
+This implies that these classes are challenging to cluster and hard to distinguish.
+While OpenScene with LSeg does not have any class with a ratio over $1$, OpenScene with OpenSeg has five: “fireplace”, “lighting”, “beam”, “lighting”, “shower”.
+These classes amount to, on average, $0.8\%$ number of voxels in the whole data set and are, therefore, rare tail classes.
+VLMaps with both encoders have a larger number of classes with a ratio over one.
+Notably, both include large classes “floor” and “table”, consisting of $4.44\%$ and $1.67\%$ of the data set, respectively.
+This implies that spatial features could be helpful in distinguishing large background classes, such as “floor”.
+VI-B2 Inter-map distinctness
+In Figure 5, the Wasserstein distances between matching, i.e., the label is the same, and non-matching sets of embeddings across all maps are shown, i.e., the blue distances show the similarity between embeddings of the same label between different maps, whereas the orange distances are similarity between embeddings of different classes.
 The medians of the data are shown with dashed lines.
+Again, a larger separation between matching and non-matching is preferable, as this would imply that classes remain distinct between environments, i.e., a door would look more like other doors than other classes.
 The ratio between the medians of matching and non-matching labels suggests that OpenScene separates the labels better than VLMaps, and LSeg separates them better than OpenSeg.
-The ratio of medians of OpenScene with LSeg is 9.21, OpenScene with OpenSeg is 3.67, VLMaps with LSeg 1.65, and VLMaps with OpenSeg 1.24, which is once again the exact ordering observed in Sections V-C1 and V-C2.
-While the shapes of distributions are similar, OpenScene has much larger distances on average, even with matching labels.
-This results from the fact that the distilled features of OpenSeg are of a larger magnitude than the image features of both VLMaps and OpenScene image features.
+The ratio of medians of OpenScene with LSeg is 6.31, OpenScene with OpenSeg is 3.78, VLMaps with LSeg 2.07, and VLMaps with OpenSeg 1.42, which is the same ordering observed in queryability results in Section VI-A.
+OpenScene has larger Wasserstein distances on average, even with matching labels, which results from the fact that the 3D features of OpenSeg are of a larger magnitude than the image features of both VLMaps and OpenScene image features.
 For example, the average norm of all OpenScene with OpenSeg embeddings in the map of the first sequence is 13.68 times larger than the average norm of all VLMaps with OpenSeg embeddings in the same map, and the maximum being 57.97 times larger than the maximum VLMaps feature.
-Furthermore, in Figure 4, the distances of matching and non-matching distances are shown per class with orange and blue bars, respectively.
+Furthermore, in Figure 6, matching and non-matching distances are shown per class with orange and blue bars, respectively.
 The same phenomena can be seen: LSeg distinguishes the classes better than OpenSeg and OpenScene better than VLMaps, which is indicated by the separation of the blue and orange box plots.
-Additionally, the ten most common and uncommon categories in the Matterport3D data set are marked with red and blue names, respectively.
+Additionally, the ten most common and uncommon categories in the Matterport3D data set that are present in each set of predictions are marked with blue and red names, respectively.
 VLMaps performs well in separating common classes and worse with uncommon classes, especially using LSeg, whereas OpenScene performs better across the whole set of labels.
-When considering all these results, the combination of OpenScene and LSeg produces maps that are both more queryable and consistent, making it the method that currently can be expected to generalize better and lead to more consistent performance in downstream robotic applications.
-VI Conclusion
-Mobile robot maps must be able to capture the semantics of the environment for the robots to perform complex tasks.
-If the semantic representation is constrained to a closed set of classes, the robot’s ability to understand the environment is constrained to the same set.
-Latent semantic maps present a way to overcome this limitation and enable open-vocabulary scene understanding by representing the semantics in a latent space with a notion of distance.
-In this work, we propose quantitative metrics to evaluate the quality of state-of-the-art latent semantic maps.
-In our experiments, OpenScene mapping using 3D scene structure with LSeg semantic embedding performed the best.
-This indicates that the 3D structure of the environment is useful for inferring the latent semantics.
+When considering all these results, the combination of OpenScene and LSeg produces maps that are more queryable and distinct, making it the method that can currently be expected to lead to more accurate performance in downstream robotic applications.
+However, OpenScene must be used with the resolution that the 3D encoder is trained on, whereas VLMaps can generalize better to different robot hardware, as it can produce maps of different resolutions, allowing the creation of maps with small memory footprint, with no significant reduction in quality.
+VII Conclusion
+In this work, we evaluate the quality of state-of-the-art latent semantic maps.
+We argue that success in downstream tasks is not alone sufficient evidence of the quality of the maps, but rather the representations themselves should be evaluated to understand their limitations.
+Therefore, we propose quantitative metrics to enable a fair comparison between methods.
+In our experiments, OpenScene mapping using 3D scene structure with LSeg semantic embedding performed the best, which indicates that the 3D structure of the environment should be used for inferring the latent semantics.
 While LSeg performed best in the semantic segmentation setting, the qualitative demonstration hints that OpenSeg might be more sensitive to rarer and more abstract queries when properly thresholded.
+Furthermore, while 2D image features are naturally scale-invariant, the 3D features are susceptible to resolution changes.
+In use cases with limited hardware resources, the image-based methods work with various resolutions, whereas 3D features must be trained for each resolution.
+This improves the adaptability of image-based methods to various use cases in real-world settings.
 The ability of visual-language maps to capture latent semantics is, however, still limited.
 Currently, fully supervised semantic segmentation methods outperform all latent semantic maps.
-The relatively low F1-scores and IoUs, especially in the instance-based test, indicate that there is still work to be done to address the problems of thresholding and clustering the query results to segment object instances.
-A data set providing ground truth open-vocabulary semantics is direly needed.
-This would allow extending the evaluation presented here to adequately address the breadth of vocabulary and capture the full promise of VLMs.
+The open-vocabulary querying method, proposed in VLMaps, comparing the query to the string “other” cannot sufficiently partition the environment, yielding poor results.
+This means that a practitioner must choose between the accuracy of the fine-tuned methods with more limited semantics and the broad vocabulary of the less accurate open-vocabulary methods.
+The relatively low F1-scores and IoUs indicate that in open-vocabulary settings, the thresholding and clustering of the query results remain an open problem.
+The growing adoption of latent semantic maps in robotics is already allowing more natural human-robot interaction and complex mobile task execution.
+The results presented in this work provide evidence that rigorously studying the qualities of these representations is important both to inform practitioners integrating these maps in their robotic pipelines and outlining open problems for researchers to explore.

### v3

@@
   4. iv)
-    We make our evaluation software publicly available [^1] The source code will be made publicly available upon acceptance of this work., enabling evaluation of the progress in this new field of research.
+    We make our evaluation software publicly available [^1] https://aalto-intelligent-robotics.github.io/latent_semantics/, enabling evaluation of the progress in this new field of research.
 II Related work
@@
   2. ii)
-    Each map $m\in\mathcal{M}$ is queried with each query $q$ in the set of queries $\mathcal{Q}$.
-The query result of a single query is a binary segmentation mask of the map, $\hat{\textbf{y}}_{q}=\{\hat{y}_{1},\dots,\hat{y}_{N}\}$, consisting of $N$ voxel masks $\hat{y}\in\{true,false\}$.
-Combined, the query results form the set of voxel-based predictions $\hat{Y}_{V}=\{\hat{\textbf{y}}_{q}|q\in\mathcal{Q}\}$.
+[…]
   3. iii)
-    For each query, the true mask $\textbf{y}_{q}$ is created from the ground truth map $m_{g}$ such, that the voxels answering query $q$ are $true$, others $false$.
+    For each query, the true mask $\textbf{y}_{q}$ is created from the ground truth map $m_{g}$ such, that the voxels answering query $q$ are truetrue, others falsefalse.
 The true masks form a set of true labels $Y_{V}=\{\textbf{y}_{q}|q\in\mathcal{Q}\}$.
@@
     The average absolute deviation $d_{l}^{i}$ is calculated for $\mathcal{E}_{l}$, and $d_{\mu}^{i}$ for $\mathcal{E}_{i}$, for each tuple $t\in\mathcal{T}$.
-The average absolute deviation $d_{a}$ measures the statistical dispersion of a set, and given a set of points =$\mathcal{E}=\{e_{1},...,e_{n}\}$ it is defined as
+The average absolute deviation $d_{a}$ measures the statistical dispersion of a set, and given a set of points $\mathcal{E}=\{e_{1},...,e_{n}\}$ it is defined as
     $d_{a}=\frac{1}{n}\sum^{n}_{j=1}|f_{d}(e_{j},\bar{\mathcal{E}})|,$ (1)
@@
 The data set was used in the original work of VLMaps, in which they selected a sample of 10 sequences from the data set.
-In this work, we use the same sequences and the same set of 40 labels [31] forming the set $\mathcal{L}$; their names are also used in the following experiments as the query set $\mathcal{Q}$ to leverage the semantic ground truth provided with the dataset [^2] Supplementary material including the list of labels will be available in arXiv and link provided here upon acceptance of this work..
+In this work, we use the same sequences and the same set of 40 labels [31] forming the set $\mathcal{L}$; their names are also used in the following experiments as the query set $\mathcal{Q}$ to leverage the semantic ground truth provided with the dataset [^2] https://aalto-intelligent-robotics.github.io/latent_semantics/.
 Since Matterport3D does not provide ground truth instance segmentation, instances used as ground truth were segmented using a region growing algorithm based on the ground truth semantics [32].
@@
 When the map is queried with a query $q$, the embeddings on the map are compared using cosine similarity to two embeddings, embedding of $q$ and embedding of string “other”.
-The binary classification mask is $true$ for voxels where the most similar query is $q$; otherwise, it is $false$.
+The binary classification mask is truetrue for voxels where the most similar query is $q$; otherwise, it is falsefalse.
 The binary mask is then post-processed by a sequence of binary closing, gaussian blurring, thresholding, and binary dilation to encompass whole object instances.
@@
 To investigate the effects of cell size on the map quality, all of the experiments were performed with a set of resolutions $r\in[0.02,0.05,0.1,0.2]$ m.
-The resolutions proposed in the original works are $0.02$m and $0.05$m in OpenScene and VLMaps, respectively.
-If not explicitly stated otherwise, we show the results in a resolution of $0.02$m for fair comparison.
+The resolutions proposed in the original works are 0.020.02m and 0.050.05m in OpenScene and VLMaps, respectively.
+If not explicitly stated otherwise, we show the results in a resolution of 0.020.02m for fair comparison.
 VI Results